# Opportunities and Challenges of Large Language Models for Low-Resource Languages in Humanities Research

### v3

@@ 5.3.2 Challenges: Limitations in Models’ Understanding of Implicit Meaning
While there are multiple innovative applications of language models in low-resource literary and religious texts, a critical challenge in their application to ancient literature and religious texts lies in their limited capacity to interpret the implicit and multilayered meanings. Such texts often draw on intricate metaphors, allegories, and symbolic references that demand not just linguistic interpretability but also cultural, historical, and philosophical insight. Existing language models, grounded in patterns derived from existing datasets, inevitably lack the depth and nuance required to faithfully interpret these complexities, and professional human translations[150] consistently outperform LLMs translations in certain contexts.
The limitation[94] is compounded by the scarcity of preserved external resources and relevant data for many low-resource languages and traditions. While some texts have survived in written form, much of the oral tradition, cultural knowledge, and historical context that breathe life into these works have been lost to time. In particular, religious and philosophical texts are highly structured performances of knowledge and authority, featuring unique rhetorical and structural patterns that are not present in everyday communication. A shallow factual summary of such texts fails to capture their underlying rhetorical and dialogic function. Without these necessary components, language models struggle to discern the deeper connections within the texts, frequently producing interpretations that are either reductive or outright erroneous. Meanwhile, LLMs often exhibit biases and cultural misrepresentations, as revealed by debate-induced evaluations, which highlight their tendency to align with dominant linguistic or cultural perspectives in training data[28], underscoring the need for fairness and contextual sensitivity in multilingual training.
@@ 6.1 Technical Challenges
LLMs face significant challenges when applied to low-resource languages due to a variety of factors, primarily stemming from the scarcity and quality of available data [67, 136, 102, 52]. Unlike high-resource languages such as English or Chinese, low-resource languages often lack the extensive and standardized datasets required for effective training. Available data is typically fragmented, non-standardized, or entirely absent, making it difficult for LLMs to learn nuanced linguistic patterns. Furthermore, critical NLP tools, such as part-of-speech taggers and annotated datasets, are frequently unavailable for these languages, further impeding the development of robust models.
Another pressing issue is the lack of computational resources in regions where low-resource languages are predominantly spoken [52, 51]. Training large-scale models demands significant computational power and infrastructure, which are often inaccessible in resource-constrained environments. The high costs associated with training and deploying these models exacerbate the problem, limiting the ability of local researchers and developers to create and utilize advanced NLP tools tailored to their linguistic and cultural needs.
The linguistic and cultural diversity inherent in low-resource languages presents additional hurdles. For example, Africa alone is home to over 2,000 languages, each with unique grammatical rules, vocabulary, and cultural contexts [131]. This diversity complicates the development of multilingual models capable of performing consistently across languages. Moreover, many low-resource languages lack sufficient digital representation, resulting in inadequate data for pretraining and instruction tuning. Consequently, LLMs trained on such data often underperform, reflecting the foundational gap in resources necessary for their development.
Low-resource languages also face specific challenges during training due to inefficiencies in tokenization processes. For instance, non-Latin scripts, such as Bengali, are often over-tokenized by standard methods like Byte Pair Encoding (BPE) [93]. Over-tokenization leads to higher computational costs and lower information density, negatively impacting model efficiency and performance. These inefficiencies highlight the need for more refined tokenization methods tailored to the characteristics of low-resource languages.
@@ 6.2 Ethical Challenges
Ethical and fairness concerns also arise when applying LLMs to low-resource languages. Models trained on limited and biased datasets may perpetuate existing linguistic or cultural biases, leading to inaccurate or inappropriate outputs. These biases can negatively impact the fair application of NLP tools in diverse cultural contexts. For instance, Low-resource Indigenous languages often encode culturally sensitive knowledge integral to their communities’ identity and heritage. Indiscriminate harvesting of Indigenous language data from online sources poses significant risks to privacy and cultural sovereignty. For example, much of the digital text available for Indigenous languages, such as South Sámi, may come from private or semi-private community contexts [138]. This makes uncritical use of such corpora problematic, as it could expose knowledge the community considers sacred or confidential. Ethical language technology development requires building trust with language communities, ensuring data use respects cultural boundaries and adheres to explicit agreements on data utilization.
LLMs, while promising, can inadvertently worsen the marginalization of low-resource languages. Low-resource Indigenous languages often have limited digitized resources, leading to minimal and potentially flawed representation in language models. Uncorrected errors in machine-generated outputs risk amplification in subsequent models, creating a feedback loop that distorts authentic language usage  [138]. Furthermore, the dominance of high-resource languages in training corpora can lead to linguistic homogenization, further weakening the already vulnerable position of low-resource languages. This imbalance threatens the linguistic diversity and the preservation of cultural identities.
@@ 7.1 Definitions and Types of Low-Resource Languages
A low-resource language refers to a language that suffers from a significant scarcity of linguistic resources, such as corpora, dictionaries, and annotated datasets. These languages often lack large-scale digital support, posing challenges for NLP and computational linguistic research [53]. Despite their diversity and cultural significance, low-resource languages are underrepresented in modern technological and computational advancements. Below, we delve into the key classifications of low-resource languages, their unique challenges, and the role of LLMs in addressing these gaps.
@@ 7.1.1 Dialects
Large language models, when augmented with transfer learning techniques or pre-trained on related high-resource languages, have the potential to bridge these gaps. However, dialects’ inherent variability still presents a major obstacle to achieving high performance in NLP applications [143].
@@ 7.1.3 Endangered Languages
Endangered languages are at risk of extinction, often spoken by small, isolated communities. Many indigenous languages in regions such as Australia, Africa, and the Americas face dwindling numbers of speakers, limited geographical reach, and minimal documentation. Cherokee language evaluation through the ChrEn dataset reveals best-case performance of 15.8/12.7 BLEU for in-domain translation, with only 2,000 fluent first-language speakers remaining. The Cherokee Nation’s loss of 346 fluent speakers between 2020-2022 underscores the urgency of developing effective NLP tools [151].
@@
  • Performance limitations: For instance, Quechua [21], even specialized models like QueEn (2024) achieve only 17.6 BLEU compared to 1.5 for standard GPT models.
@@ 7.2 Current State of Foundation Models for Low-Resource Languages
As of August 2025, the landscape of LLMs has evolved significantly with the confirmed releases of GPT-5 (OpenAI, August 7, 2025), Gemini 2.5 Pro (Google, March-July 2025), and Claude 4 (Anthropic, May 2025). GPT-5 shows performance improvements with 94.6% accuracy on AIME 2025 and 74.9% on SWE-bench Verified, featuring unified reasoning architecture and approximately 45% fewer factual errors than GPT-4o [105]. Gemini 2.5 Pro introduces adaptive reasoning budgets with “thinking” capabilities and native multimodality across text, images, audio, and video with 1M+ token context windows [27]. Claude Sonnet 4 offers hybrid reasoning models with 200K context windows standard and 1M tokens available in beta, achieving 96% accuracy across 12 major languages [7].
Despite these advances, a fundamental disparity persists between high-resource and low-resource language performance. The Stanford HAI White Paper [109] identifies that ”most major LLMs underperform for non-English—and especially low-resource—languages; are not attuned to relevant cultural contexts; and are not accessible in parts of the Global South”. This gap becomes particularly pronounced when examining specific low-resource languages, where even state-of-the-art models frequently perform below random baselines.
@@ 7.3 Scarcity and Complexity of Corpora in Low-Resource Languages
Low-resource languages are characterized by the scarcity of corpora, which significantly hinders their integration into modern NLP systems [143]. Unlike high-resource languages such as English and French, which have benefited from decades of extensive corpus development, these languages often lack sufficient datasets. The challenges in data collection are multifaceted, involving logistical difficulties, financial constraints, and cultural sensitivities. Assembling even basic corpora for such languages can be a significant undertaking, and the lack of resources limits the development of NLP models tailored to these linguistic contexts.
The available datasets for low-resource languages often suffer from poor quality and limited scope, making them suboptimal for robust model training [115]. Inconsistencies within the datasets and the absence of comprehensive linguistic coverage further restrict their utility. The MiLiC-Eval  [147] benchmark reveals that Tibetan requires 70-430 tokens per sentence compared to 27 for English, highlighting fundamental architectural limitations in handling morphologically complex scripts.
@@
The linguistic complexity of low-resource languages further amplifies these challenges. Many of these languages possess unique grammatical structures, such as the lack of clear distinctions between grammatical categories or the use of non-linear syntax [115]. These features make it difficult to adapt existing NLP techniques designed for high-resource languages. Moreover, orthographic diversity adds another layer of difficulty. Some languages lack standardized writing systems, while others use scripts that are challenging to digitize or process with current computational tools.
Additionally, regional variations within low-resource languages pose a significant obstacle to corpus development. Differences in pronunciation, vocabulary, and syntax across geographical areas require highly localized datasets to capture the full linguistic diversity of these languages [143]. Such regional variations demand substantial effort to develop representative and contextually accurate corpora. These factors collectively underscore the pressing need for innovative strategies to overcome the intertwined issues of scarcity and complexity in low-resource language processing.
@@ 7.4 Role of Large Language Models in Addressing Challenges
Zhang et al. established fundamental principles and proposing machine-in-the-loop processing approaches that prioritize community engagement  [151]. This work demonstrates that successful low-resource language NLP requires not merely technical innovation but collaborative frameworks that respect indigenous data sovereignty and cultural contexts. The TLUE benchmark for Tibetan  [42], which reveals that most current LLMs including GPT-4 perform below the 25% random baseline on Tibetan tasks, with Claude-3.5-Sonnet achieving the highest accuracy at only 35.6%. Similarly, the AmericasNLI dataset  [34] covering ten Indigenous languages of the Americas shows that XLM-R’s zero-shot performance averages merely 38.48% accuracy, only slightly above random chance.
@@ 7.5.1 Transfer Learning
Transfer learning is a widely adopted technique to address the challenges posed by low-resource languages [158]. By pre-training a model on high-resource languages and fine-tuning it on low-resource languages, the model can acquire foundational linguistic knowledge from the former and adapt to the unique characteristics of the latter.
Recent advances in chain-of-thought reasoning have revolutionized transfer learning for low-resource languages. Research demonstrates that scaling inference-time compute through CoT reasoning enables models to leverage high-resource language knowledge more effectively [144]. Fine-tuning on just 1,000 reasoning traces in languages like Swahili produces 33.8% performance gains, as the model learns to use well-represented languages as internal pivot points during reasoning. This represents a paradigm shift from static knowledge transfer to dynamic, reasoning-based transfer where the model explicitly generates intermediate steps in high-resource languages before producing low-resource language outputs.
@@ 7.5.2 Cross-Language Pretraining
Cross-language pretraining has evolved from early multilingual models to sophisticated architectures that better handle linguistic diversity. Initial approaches, exemplified by models like mBERT, XLM, and XLM-R [46], train on multilingual corpora to learn shared representations across languages.
@@
However, a fundamental limitation of dense multilingual models is parameter interference—all languages compete for the same parameter space, often resulting in high-resource languages dominating the representations. The emergence of sparse Mixture of Experts (MoE) architectures directly addresses this challenge [157]. Unlike dense models, MoE maintains dedicated expert networks for different language families, preventing parameter interference while preserving the benefits of shared learning.
@@ 7.5.3 Multi-Task Learning
The introduction of dynamic routing mechanisms, particularly in GPT-5’s unified architecture, has elevated these benefits to new levels [105]. Rather than treating all tasks equally, the model now employs a real-time router that dynamically allocates computational resources based on task complexity and linguistic demands. For low-resource languages, this means:
@@ 7.5.4 Data Augmentation
Data augmentation addresses the problem of insufficient training data by generating additional examples and leveraging available resources more effectively [114]. This technique enriches datasets through both preprocessing and inference-time strategies, providing models with more diverse and representative samples.
@@
  • Dynamic resource integration: LLMs like GPT-5 and Claude 4 can alternate between reasoning and tool use during inference, incorporating external resources such as dictionaries and parallel corpora as needed [105, 7]. This enables real-time augmentation where the model searches for and integrates relevant linguistic resources based on specific query requirements.
Expanded context windows have fundamentally changed how augmentation can be applied. With GPT-5 supporting 400K tokens and Claude Sonnet 4 supporting 1M tokens [105, 7], entire low-resource language corpora can be provided as context during inference. This capability enables several key advantages: First, for many endangered languages, the complete written corpus fits within a single context window, effectively providing the model with all available knowledge at inference time. Second, models can become temporarily specialized for specific languages by loading relevant resources into context, eliminating the need for separate fine-tuning. Third, historical texts and oral tradition transcriptions can be used directly without risk of corruption through synthetic generation processes.

### v4

@@ 5.3.2 Challenges: Limitations in Models’ Understanding of Implicit Meaning
While there are multiple innovative applications of language models in low-resource literary and religious texts, a critical challenge in their application to ancient literature and religious texts lies in their limited capacity to interpret the implicit and multilayered meanings. Such texts often draw on intricate metaphors, allegories, and symbolic references that demand not just linguistic interpretability but also cultural, historical, and philosophical insight. Existing language models, grounded in patterns derived from existing datasets, inevitably lack the depth and nuance required to faithfully interpret these complexities, and professional human translations[zhang2024goodllmsliterarytranslation] consistently outperform LLMs translations in certain contexts.
The limitation[Mahowald_2024] is compounded by the scarcity of preserved external resources and relevant data for many low-resource languages and traditions. While some texts have survived in written form, much of the oral tradition, cultural knowledge, and historical context that breathe life into these works have been lost to time. In particular, religious and philosophical texts are highly structured performances of knowledge and authority, featuring unique rhetorical and structural patterns that are not present in everyday communication. A shallow factual summary of such texts fails to capture their underlying rhetorical and dialogic function. Without these necessary components, language models struggle to discern the deeper connections within the texts, frequently producing interpretations that are either reductive or outright erroneous. Meanwhile, LLMs often exhibit biases and cultural misrepresentations, as revealed by debate-induced evaluations, which highlight their tendency to align with dominant linguistic or cultural perspectives in training data[demidova-etal-2024-john], underscoring the need for fairness and contextual sensitivity in multilingual training.
@@ 6.1 Technical Challenges
LLMs face significant challenges when applied to low-resource languages due to a variety of factors, primarily stemming from the scarcity and quality of available data [kholodna2024llms, wang2024improving, ogueji2021small, hedderich2020survey]. Unlike high-resource languages such as English or Chinese, low-resource languages often lack the extensive and standardized datasets required for effective training. Available data is typically fragmented, non-standardized, or entirely absent, making it difficult for LLMs to learn nuanced linguistic patterns. Furthermore, critical NLP tools, such as part-of-speech taggers and annotated datasets, are frequently unavailable for these languages, further impeding the development of robust models.
Another pressing issue is the lack of computational resources in regions where low-resource languages are predominantly spoken [hedderich2020survey, hasan2024large]. Training large-scale models demands significant computational power and infrastructure, which are often inaccessible in resource-constrained environments. The high costs associated with training and deploying these models exacerbate the problem, limiting the ability of local researchers and developers to create and utilize advanced NLP tools tailored to their linguistic and cultural needs.
The linguistic and cultural diversity inherent in low-resource languages presents additional hurdles. For example, Africa alone is home to over 2,000 languages, each with unique grammatical rules, vocabulary, and cultural contexts [tonja2024inkubalm]. This diversity complicates the development of multilingual models capable of performing consistently across languages. Moreover, many low-resource languages lack sufficient digital representation, resulting in inadequate data for pretraining and instruction tuning. Consequently, LLMs trained on such data often underperform, reflecting the foundational gap in resources necessary for their development.
Low-resource languages also face specific challenges during training due to inefficiencies in tokenization processes. For instance, non-Latin scripts, such as Bengali, are often over-tokenized by standard methods like Byte Pair Encoding (BPE) [mahfuz2024too]. Over-tokenization leads to higher computational costs and lower information density, negatively impacting model efficiency and performance. These inefficiencies highlight the need for more refined tokenization methods tailored to the characteristics of low-resource languages.
@@ 6.2 Ethical Challenges
Ethical and fairness concerns also arise when applying LLMs to low-resource languages. Models trained on limited and biased datasets may perpetuate existing linguistic or cultural biases, leading to inaccurate or inappropriate outputs. These biases can negatively impact the fair application of NLP tools in diverse cultural contexts. For instance, Low-resource Indigenous languages often encode culturally sensitive knowledge integral to their communities’ identity and heritage. Indiscriminate harvesting of Indigenous language data from online sources poses significant risks to privacy and cultural sovereignty. For example, much of the digital text available for Indigenous languages, such as South Sámi, may come from private or semi-private community contexts [wiechetek2024ethical]. This makes uncritical use of such corpora problematic, as it could expose knowledge the community considers sacred or confidential. Ethical language technology development requires building trust with language communities, ensuring data use respects cultural boundaries and adheres to explicit agreements on data utilization.
LLMs, while promising, can inadvertently worsen the marginalization of low-resource languages. Low-resource Indigenous languages often have limited digitized resources, leading to minimal and potentially flawed representation in language models. Uncorrected errors in machine-generated outputs risk amplification in subsequent models, creating a feedback loop that distorts authentic language usage  [wiechetek2024ethical]. Furthermore, the dominance of high-resource languages in training corpora can lead to linguistic homogenization, further weakening the already vulnerable position of low-resource languages. This imbalance threatens the linguistic diversity and the preservation of cultural identities.
@@ 7.1 Definitions and Types of Low-Resource Languages
A low-resource language refers to a language that suffers from a significant scarcity of linguistic resources, such as corpora, dictionaries, and annotated datasets. These languages often lack large-scale digital support, posing challenges for NLP and computational linguistic research [Hedderich2020ASO]. Despite their diversity and cultural significance, low-resource languages are underrepresented in modern technological and computational advancements. Below, we delve into the key classifications of low-resource languages, their unique challenges, and the role of LLMs in addressing these gaps.
@@ 7.1.1 Dialects
Large language models, when augmented with transfer learning techniques or pre-trained on related high-resource languages, have the potential to bridge these gaps. However, dialects’ inherent variability still presents a major obstacle to achieving high performance in NLP applications [LSSZ202408005].
@@ 7.1.3 Endangered Languages
Endangered languages are at risk of extinction, often spoken by small, isolated communities. Many indigenous languages in regions such as Australia, Africa, and the Americas face dwindling numbers of speakers, limited geographical reach, and minimal documentation. Cherokee language evaluation through the ChrEn dataset reveals best-case performance of 15.8/12.7 BLEU for in-domain translation, with only 2,000 fluent first-language speakers remaining. The Cherokee Nation’s loss of 346 fluent speakers between 2020-2022 underscores the urgency of developing effective NLP tools [zhang2022cherokee].
@@
  • Performance limitations: For instance, Quechua [chen2024queen], even specialized models like QueEn (2024) achieve only 17.6 BLEU compared to 1.5 for standard GPT models.
@@ 7.2 Current State of Foundation Models for Low-Resource Languages
As of August 2025, the landscape of LLMs has evolved significantly with the confirmed releases of GPT-5 (OpenAI, August 7, 2025), Gemini 2.5 Pro (Google, March-July 2025), and Claude 4 (Anthropic, May 2025). GPT-5 shows performance improvements with 94.6% accuracy on AIME 2025 and 74.9% on SWE-bench Verified, featuring unified reasoning architecture and approximately 45% fewer factual errors than GPT-4o [openai2025gpt5]. Gemini 2.5 Pro introduces adaptive reasoning budgets with “thinking” capabilities and native multimodality across text, images, audio, and video with 1M+ token context windows [google2025gemini]. Claude Sonnet 4 offers hybrid reasoning models with 200K context windows standard and 1M tokens available in beta, achieving 96% accuracy across 12 major languages [anthropic2025claude4].
Despite these advances, a fundamental disparity persists between high-resource and low-resource language performance. The Stanford HAI White Paper [stanford2025hai] identifies that ”most major LLMs underperform for non-English—and especially low-resource—languages; are not attuned to relevant cultural contexts; and are not accessible in parts of the Global South”. This gap becomes particularly pronounced when examining specific low-resource languages, where even state-of-the-art models frequently perform below random baselines.
@@ 7.3 Scarcity and Complexity of Corpora in Low-Resource Languages
Low-resource languages are characterized by the scarcity of corpora, which significantly hinders their integration into modern NLP systems [LSSZ202408005]. Unlike high-resource languages such as English and French, which have benefited from decades of extensive corpus development, these languages often lack sufficient datasets. The challenges in data collection are multifaceted, involving logistical difficulties, financial constraints, and cultural sensitivities. Assembling even basic corpora for such languages can be a significant undertaking, and the lack of resources limits the development of NLP models tailored to these linguistic contexts.
The available datasets for low-resource languages often suffer from poor quality and limited scope, making them suboptimal for robust model training [ranathunga2023neural]. Inconsistencies within the datasets and the absence of comprehensive linguistic coverage further restrict their utility. The MiLiC-Eval  [zhang2025milic] benchmark reveals that Tibetan requires 70-430 tokens per sentence compared to 27 for English, highlighting fundamental architectural limitations in handling morphologically complex scripts.
@@
The linguistic complexity of low-resource languages further amplifies these challenges. Many of these languages possess unique grammatical structures, such as the lack of clear distinctions between grammatical categories or the use of non-linear syntax [ranathunga2023neural]. These features make it difficult to adapt existing NLP techniques designed for high-resource languages. Moreover, orthographic diversity adds another layer of difficulty. Some languages lack standardized writing systems, while others use scripts that are challenging to digitize or process with current computational tools.
Additionally, regional variations within low-resource languages pose a significant obstacle to corpus development. Differences in pronunciation, vocabulary, and syntax across geographical areas require highly localized datasets to capture the full linguistic diversity of these languages [LSSZ202408005]. Such regional variations demand substantial effort to develop representative and contextually accurate corpora. These factors collectively underscore the pressing need for innovative strategies to overcome the intertwined issues of scarcity and complexity in low-resource language processing.
@@ 7.4 Role of Large Language Models in Addressing Challenges
Zhang et al. established fundamental principles and proposing machine-in-the-loop processing approaches that prioritize community engagement  [zhang2022cherokee]. This work demonstrates that successful low-resource language NLP requires not merely technical innovation but collaborative frameworks that respect indigenous data sovereignty and cultural contexts. The TLUE benchmark for Tibetan  [gao2025tlue], which reveals that most current LLMs including GPT-4 perform below the 25% random baseline on Tibetan tasks, with Claude-3.5-Sonnet achieving the highest accuracy at only 35.6%. Similarly, the AmericasNLI dataset  [ebrahimi-2022-americasnli] covering ten Indigenous languages of the Americas shows that XLM-R’s zero-shot performance averages merely 38.48% accuracy, only slightly above random chance.
@@ 7.5.1 Transfer Learning
Transfer learning is a widely adopted technique to address the challenges posed by low-resource languages [zoph2016transfer]. By pre-training a model on high-resource languages and fine-tuning it on low-resource languages, the model can acquire foundational linguistic knowledge from the former and adapt to the unique characteristics of the latter.
Recent advances in chain-of-thought reasoning have revolutionized transfer learning for low-resource languages. Research demonstrates that scaling inference-time compute through CoT reasoning enables models to leverage high-resource language knowledge more effectively [yeo2025longcot]. Fine-tuning on just 1,000 reasoning traces in languages like Swahili produces 33.8% performance gains, as the model learns to use well-represented languages as internal pivot points during reasoning. This represents a paradigm shift from static knowledge transfer to dynamic, reasoning-based transfer where the model explicitly generates intermediate steps in high-resource languages before producing low-resource language outputs.
@@ 7.5.2 Cross-Language Pretraining
Cross-language pretraining has evolved from early multilingual models to sophisticated architectures that better handle linguistic diversity. Initial approaches, exemplified by models like mBERT, XLM, and XLM-R [gonen2020s], train on multilingual corpora to learn shared representations across languages.
@@
However, a fundamental limitation of dense multilingual models is parameter interference—all languages compete for the same parameter space, often resulting in high-resource languages dominating the representations. The emergence of sparse Mixture of Experts (MoE) architectures directly addresses this challenge [zhou2022moe]. Unlike dense models, MoE maintains dedicated expert networks for different language families, preventing parameter interference while preserving the benefits of shared learning.
@@ 7.5.3 Multi-Task Learning
The introduction of dynamic routing mechanisms, particularly in GPT-5’s unified architecture, has elevated these benefits to new levels [openai2025gpt5]. Rather than treating all tasks equally, the model now employs a real-time router that dynamically allocates computational resources based on task complexity and linguistic demands. For low-resource languages, this means:
@@ 7.5.4 Data Augmentation
Data augmentation addresses the problem of insufficient training data by generating additional examples and leveraging available resources more effectively [ragni2014data]. This technique enriches datasets through both preprocessing and inference-time strategies, providing models with more diverse and representative samples.
@@
  • Dynamic resource integration: LLMs like GPT-5 and Claude 4 can alternate between reasoning and tool use during inference, incorporating external resources such as dictionaries and parallel corpora as needed [openai2025gpt5, anthropic2025claude4]. This enables real-time augmentation where the model searches for and integrates relevant linguistic resources based on specific query requirements.
Expanded context windows have fundamentally changed how augmentation can be applied. With GPT-5 supporting 400K tokens and Claude Sonnet 4 supporting 1M tokens [openai2025gpt5, anthropic2025claude4], entire low-resource language corpora can be provided as context during inference. This capability enables several key advantages: First, for many endangered languages, the complete written corpus fits within a single context window, effectively providing the model with all available knowledge at inference time. Second, models can become temporarily specialized for specific languages by loading relevant resources into context, eliminating the need for separate fine-tuning. Third, historical texts and oral tradition transcriptions can be used directly without risk of corruption through synthetic generation processes.